Network engineers must understand and apply different number bases, especially when designing networks and subnets. For this discussion question, select one of the following problems and post the solution showing a step-by-step solution. Provide an example of how you could apply your solution into a real-world technology or similar context.

  1. Convert 1000100011111 from base 2 to base 10.
  2. Convert 28EAB from Base 16 to base 10.
  3. Convert 6758 from base 10 to base 2.
  4. Convert 88898 from base 10 to base 16.
  5. Convert CDDEA from base 16 to base 2.
  6. Convert 11101010111001 from base 2 to base 16.
  7. Convert 55786 from base 10 to 8.
  8. Convert EDDEAA from base 16 to 8.
  9. Convert 10000100111 from base 2 to base 10.
  10. Convert 9234 from base 10 to base 2.
Answer #1

Convert 1000100011111 from base 2 to base 10.

1000100011111 = 1*212 + 0*211 + 0*210 + 0*29 + 1*28 + 0*27 + 0*26 + 0*25 + 1*24 + 1*23 + 1*22 + 1*21 + 1*20

= 4096 + 256 + 16 + 8 + 4 + 2 + 1 = 4383

Convert 28EAB from Base 16 to base 10.

28EAB = 2*164 + 8*163​​​​​​​ + 14*162 + 10*161 + 11*160​​​​​​ = 131072 + 32768 + 3584 + 160 + 11 = 167595

Convert 6758 from base 10 to base 2.

6758 = 11010011001102

Convert 88898 from base 10 to base 16.

88898 = 15B4216

NOTE: As per Chegg policy, I am allowed to answer only 4 questions (including sub-parts) on a single post. Kindly post the remaining questions separately and I will try to answer them. Sorry for the inconvenience caused.